OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 - 9x + 18

OpenStudy (maheshmeghwal9):

HEY \[\LARGE{\color{red}{TRINOMIAL?}}\]

OpenStudy (anonymous):

x^2 - 9x + 18 = x^2 -3x - 6x +18 = x ( x-3) - 6 (x-3) = ... ?

OpenStudy (maheshmeghwal9):

it should be binomial becoz it has a power two

OpenStudy (anonymous):

It's a trinomial because there are 3 terms

OpenStudy (anonymous):

Please see this: http://www.mathwords.com/t/trinomial.htm